Wednesday 20 October 1999 11.00am EST
|During the last 24 hours Chandra completed observations of the GTO targets Cygnus X-1 and Cygnus X-3 with ACIS-S and HETG. During the next 24 hours observations will be conducted of PKS 2155-304 and 4U1636-53 with ACIS-S and the HETG, and an observation started of the target A697 with ACIS-I.
The re-built load containing re-configured ACIS parameter blocks to avoid use of FEP-0 was uplinked during the evening of 10/19/99 and is operating as expected. This load will continue through 12AM 10/24 when the next mission schedule will start. The products for this schedule are expected to be in review today.
The spacecraft systems continued to operate nominally.
Return to Status Reports